# Fisher information, Crámer-Rao complexity and temperature in logistic map

### v1

@@ Abstract
In this work we study the dynamics of the logistic map based on a probabilistic characterization in terms of the invariant density. We analyze the relevant regimes of the dynamics (regular, oscillatory, onset chaotic and fully chaotic) in terms of the Fisher information and the Crámer-Rao (CR) complexity. We found that these informational quantifiers allow to distinguish the dynamical regions of the map, by maximizing the Fisher information in the regular behavior and with the CR complexity exhibiting variations and a maximum near to the Pomeau-Maneville scenario. Fisher information as a function of time is examined to the light of the Frieden’s informational interpretation of the Thermodynamics Second Law. Finally, we apply the Equipartition Theorem to propose a definition of temperature for the logistic map, thus providing a macroscopic signature of the dynamics.
@@ I Introduction
Maps constitute useful tools to study the discrete evolution of a system expressed by $x_{i+1}=f(x_{i})$, with $f(x)$ a function describing the dynamics. The logistic map constitutes one of the most interesting objects due to its several applications in physics, biology, chemistry, engineering science and mathematics [1]. The iteration of its quadratic formula reveals many types of dynamical regimes going from the regular ones to the fully chaotic one, and between them exhibiting intermittency, aperiodicity, and regular-chaotic mixed behaviors, thus manifesting the richness of the logistic map for testing chaos, ergodicity, onset of chaos, complexity and serving for population growth approaches, among others. Complementarily, a probabilistic description for discrete systems is the given by the theory of the invariant measures and Markov operators [2]. In this approach, the features of the dynamics are studied from the evolution of probabilistic densities that characterize the knowledge of the system at a given instant. The evolution of any density is obtained from the iterations of the Frobenius-Perron operator over an arbitrary initial density. Of particular importance is the existence of invariant densities of the Frobenius-Perron operator since they represent stationary states modeling the system in the asymptotic limit of large times.
@@
On the other hand, statistical quantifiers have been shown to be important theoretical tools for characterizing several phenomena. Between these ones, the Fisher information (FI) has attracted special attention in the literature. FI arose from the problem of parameter estimation in the context of statistics, and since then it has been employed for characterizing pattern genetics in biology, dynamical phase transitions in environmental engineering, chemistry, economy, quantum mechanics, deduction of motion equations, by making use of its sensibility against local variations of the probability distribution. Intuitively, the Fisher information can distinguish situations with behaviors being radically different. For instance, a regular behavior characterized by a piked distribution gives a high valued Fisher information, while a fully chaotic dynamics represented by a uniform distribution minimizes the Fisher information. The CR complexity plays a complementary role by distinguishing behaviors where the Fisher information is not able for. When we have different dynamical regimes with similar values of the Fisher information, we can calculate the CR complexity for distinguishing between them. Additionally, FI has been studied to give an alternative formulation of the Second Law of Thermodynamics. Illustrating with several examples and theoretical arguments, in [3] Frieden has suggested that FI can provide a interpretation of the arrow of time, by postulating that Second Law of Thermodynamics is equivalent to the condition $dI/dt\leq 0$. Physically, this inequality means that if the FI level of a system is observed to be decreasing, that is $dI<0$, then its history is necessarily advancing, which is the main content of the “$I$-Theorem” (Section 1.8.2 of [3]).
@@
The main goal of this work is to characterize some relevant regimes of the logistic map dynamics from a probabilistic description in terms of the invariant density, from which we calculate the Fisher information (FI) and its associated Cramér-Rao (CR) complexity for analyzing dynamical transitions. In addition, we propose a notion of temperature based on the Equipartition Theorem, which allows to provide a macroscopic signature of the map along with a study of the behavior of the Thermodynamics Second Law by means of the Frieden’s informational interpretation. The work is organized as follows.
@@ II.1 Probability density functions, discrete maps and invariant density
A probability density function defined over an abstract space $\Gamma$ (typically, a subset of $\mathbb{R}^{m}$) is any nonnegative function $p:\Gamma\rightarrow\mathbb{R}_{+}$ such that $\int_{\Gamma}p(x)dx=1$ with $\Gamma$ called the space of events. In an experiment, $\Gamma$ represents the set of all possible outcomes. If these ones are $W$ possible results, then one has a discrete probability distribution expressed by a probability vector $(p_{1},p_{2},\ldots,p_{W})$ being $p_{i}$ the probability that the ith result occurs and the normalization condition reads $\sum_{i=1}^{W}p_{i}=1$.
Given a set $\Gamma$ and a continuous function $f:\Gamma\rightarrow\Gamma$, we say that the sequence $\{x_{n}\}_{n\in\mathbb{N}_{0}}\subseteq\Gamma$ such that
@@
defines a discrete map. Given $x_{0}\in\Gamma$ as the starting point, the sets $\{x_{n}:n\in\mathbb{N}_{0}\}$ are called orbits, where $x_{\infty}=\lim_{n\rightarrow\infty}x_{n}$ stands for the limit of a orbit. From the physical viewpoint, a discrete map models a system whose dynamics is given by iterating (1), where each iteration corresponds to a time step. If the system is initially in a state $x_{0}$, then $x_{n}$ represents the state after $n$ time steps. When there exists an element $c\in\Gamma$ such that $f(x^{*})=x^{*}$, we say that $x^{*}$ is a fixed point of $f$ and they are physically interpreted as stationary states of the system. The dynamics of a discrete map can be characterized in terms of probability density functions as follows. The Frobenius–Perron operator $P:\mathbb{L}^{1}(\Gamma)\rightarrow\mathbb{L}^{1}(\Gamma)$ associated with the map (1) is given by [2]
@@
is called the invariant density of the map, constituting a invariant measure. Numerically, the invariant density $\rho(x)$ can be constructed by dividing $\Gamma$ into $W$ intervals and defining
@@
with $1_{[\frac{i-1}{W},\frac{i}{W})}(x)$ standing for the characteristic function of the interval $\frac{i-1}{W},\frac{i}{W})$. It is worth to be noted that in the limit $N\rightarrow\infty$ the invariant density $\rho(x)$ does not depend on the starting point $x_{0}$.
@@ II.2 Logistic map
where $\mu\in(0,4]$ is the external parameter and $x_{0}\in\Gamma=[0,1]$ the initial condition. The case $\mu=0$ is not of interest since we have $x_{n+1}=0$ for all $n\in\mathbb{N}_{0}$. We review some relevant regimes that we will employ for characterizing them from the formalism presented. When $0<\mu\leq 3$ all the orbits are convergent and we have that $x_{\infty}=0$ for $0<\mu<1$ and $x_{\infty}=\frac{\mu-1}{\mu}$ for $1\leq\mu\leq 3$ with different rates of convergence. For $0<\mu<3$ the orbits rapidly approach to the asymptotic limit with a linear convergence for $2\leq\mu<3$ and less than linear for $\mu=3$. When $3\leq\mu\leq 3.44949$ the orbits oscillate between two values and with $3.44949<\mu\leq 3.56995$ almost all the orbits oscillate between four values. The value $\mu=3.56995$ represents the beginning of the chaotic behavior (onset of chaos) where orbit oscillations of finite period are not observed. The region $3.56995\leq\mu\leq 3.82843$ represents the called Pameau-Maneville scenario, with the orbits manifesting a periodic laminar phase interrupted by bursts of aperiodic behavior. For $\mu=4$ the fully chaotic behavior emerges with all the orbits dense in $[0,1]$ provided with a mixing dynamics. Interestingly, when $\mu=2$ and $\mu=4$ there exist exact solutions for the orbits, given by
@@ II.3 Fisher information and Crámer-Rao complexity
The discrete version of the Fisher (FI) information can be given in terms of the expression
@@
for all discrete probability density $\{p_{i}\}$, $p_{i}\geq 0$ and $\sum p_{i}=1$. The associated statistical quantifier to the Fisher information is the Crámer-Rao complexity, given by
@@
where $\sigma^{2}$ stands for the variance of the probability density $\{p_{i}\}$, that is
@@
and $z_{i}\in[\frac{i-1}{W},\frac{i}{W})$ with $p=(p_{1},p_{2},\ldots,p_{W})$ a discrete probability density. The Crámer-Rao bound states that the CR complexity is low bounded by 1, i.e. $C[p]\geq 1$, being maximized for the Gaussian probability densities. We only focus on the Fisher information and the CR complexity so the employment of other measures of information and its associated complexities [4] is out of the scope of this work.
@@ II.4 Equipartition Theorem
Equipartition Theorem (ET) establishes a relation between a temperature of a system and the average of the total kinetic energy of their particles, expressed in the general form by
@@
where the brackets means an ensemble average provided with the Liouville equilibrium density in the asymptotic limit $t\rightarrow\infty$, $H$ is the Hamiltonian of the system, $k_{B}$ is the Boltzmann constant, $T$ is the temperature and $q_{n}$ is the nth coordinate of the system in phase space. For the special case of a system composed by non-interacting particles of mass $m$, the average energy per particle $E_{\textrm{particle}}=\langle H_{\textrm{particle}}\rangle$ is given by the Hamiltonian $H_{particle}=(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})/2m$, so applying (13) for $q_{i}=p_{i}$ ($i=x,y,z$) it follows
@@
from which we can case that temperature of the system is proportional to the average energy per particle. It is worth noting that some hypotheses are needed for the validity of ET. Typically, ET holds for ergodic systems in thermal equilibrium with a density distribution in phase space equiprobable for all the states. For our proposal it is convenient to make (14) dimensionless, so considering the one-dimensional case, from (14) we can deduce the dimensionless temperature $T/T_{0}$ as
@@
where $T_{0}$ and $v_{0}$ are arbitrary characteristic temperature and velocity for the system.
@@ III Statistical quantifiers and temperature for logistic map
We calculate the Fisher information and its CR complexity for a set of values of the parameter $\mu$ of the logistic map. Next, for some representative values of $\mu$, we analyze the temporal evolution of the Fisher information from the Frieden’s interpretation of Second Law of Thermodynamics. Finally, by employing the Equipartition Theorem an associated temperature for the logistic map is obtained along with the FI and the CR complexity as a function of the temperature.
III.1 Fisher information and CR complexity of logistic map
For the numerical calculations of the FI and the CR complexity we employed $N=10^{6}$ steps and $W=10^{4}$ bins, thus guaranteeing $N\gg W$. By increasing $N$ the results remained the same due to the convergence of the invariant density. In figures 1 and 2 the FI and the CR complexity in function of the parameter $\mu$ are illustrated.
@@ III.1 Fisher information and CR complexity of logistic map
From Fig. 2 we see that FI presents an absolute maximum value for $\mu=1$, which corresponds to regular behavior with orbits converging to $x_{\infty}=0$. For $0\leq\mu<1$ and $1\leq\mu<3.4$ the FI behaves approximately constant, from which it is not possible to distinguish the dynamical transitions for these values. From $\mu=3.4$ the FI decreases up to the region $3.8<\mu<3.9$ where a local maximum occurs. FI in the fully chaotic regime $\mu=4$ has the same value as for $3.7\leq\mu\leq 3.8$. By contrast, from Fig. 2 we see that the CR complexity provides a distinction between
In Fig. 3 we see how the CR complexity takes high values while the FI is vanishingly small, and vice-versa.
@@
III.2 Evolution of Fisher information and CR complexity of logistic map
With the aim of testing the Frieden’s informational interpretation of the Second Law we analyze the FI and the CR complexity as functions of time.
III.3 Temperature associated to logistic map
Considering that $v_{i}=x_{i+i}-x_{i}$ represents a dimensionless average velocity of the logistic map between the ith and (i+1)th steps, then we apply the relationship (15) with $mv_{0}^{2}/(k_{B}T_{0})=1$ and propose the map temperature (MP) at Nth step, denoted by $T(\mu,N,M)$ for the logistic map as follows:

### v2

# Invariant density statistical quantifiers and a temperature for the logistic map

@@ Abstract
In this work, we study the dynamics of the logistic map based on a probabilistic characterization in terms of the invariant density. We analyze the relevant regimes of the dynamics (regular, oscillatory, onset chaotic and fully chaotic) in terms of the Fisher information and the Crámer-Rao (CR) complexity. We have found that these informational quantifiers allow to distinguish the dynamical regions of the map, by maximizing the Fisher information in the regular behavior and with the CR complexity exhibiting variations and a maximum near to the Pameau-Maneville scenario. Fisher information as a function of time is examined in the light of Frieden’s informational interpretation of the Second Law of Thermodynamics. We apply the Equipartition Theorem to propose a definition of temperature for the logistic map, providing a macroscopic signature of the dynamics.
@@ I Introduction
Discrete maps of the form $x_{i+1}=f(x_{i})$ constitute a powerful tool in the study of dynamical systems. Among those, the logistic map stands out as one of the most interesting ones, possessing several applications in physics, biology, chemistry, engineering, econophysics, mathematics and science in general [1, 2, 3, 4]. The iteration of its simple quadratic formula reveals an unexpected richness allowing to describe distinct dynamical regimes going from regular to fully chaotic ones, passing in between through intermittency, aperiodicity and regular-chaotic mixed behaviors, manifesting its ability for testing chaos, ergodicity, onset of chaos, complexity and serving as building block for more complex model constructions regarding population growth, epidemic control, quantum systems and many others [5, 6, 7, 8]. Complementarily, a probabilistic description for discrete systems is given by the theory of invariant measures and Markov operators [9, 10, 11]. In that approach, the dynamical system features are studied from the probability densities evolution characterizing the knowledge of the system at a given instant. The evolution of any density is obtained from the iterations of the Frobenius-Perron operator over an arbitrary initial density. Of particular importance is the existence of invariant densities of the Frobenius-Perron operator since they represent stationary states modeling the system in the asymptotic limit of large times.
On the other hand, statistical quantifiers have been shown to be important theoretical tools for characterizing several phenomena in physics and related areas. Notably, the Fisher information (FI) concept [12, 13] can be employed for exploring several phenomena like statistical ensembles, environmental systems, quantum states, information in inhomogeneous systems, etc [14, 15, 16, 17]. FI arose from the problem of parameter estimation in the context of statistics, and since then it has been employed for characterizing pattern genetics in biology, dynamical phase transitions in environmental engineering, chemistry, economy, quantum mechanics, deduction of motion equations, by making use of its sensibility against local variations of the probability distribution. For a given one-dimensional continuous probability density function $p(s)$, the corresponding Fisher information $I$ can be defined as (Eq. (1.2) of [13])
$I=\int dx\,{p^{\prime}(x)}^{2}/p(x)\,.$ (1)
@@
Intuitively, FI can distinguish situations with behaviors being radically different. For instance, a regular behavior characterized by a piked distribution gives a high value for $I$, while a fully chaotic dynamics represented by a uniform distribution minimizes FI. The Crámer-Rao (CR) complexity plays a complementary role by distinguishing behaviors where FI alone is not sufficient. When we have different dynamical regimes with similar values for the Fisher information, we can calculate the CR complexity for distinguishing between them. Additionally, FI has been studied to give an alternative formulation of the Second Law of Thermodynamics. Illustrating with several examples and theoretical arguments, in [13], Frieden has suggested that FI can provide a interpretation of the arrow of time, by postulating that Second Law of Thermodynamics is equivalent to the condition $dI/dt\leq 0$. Physically, this inequality means that if the FI level of a system is observed to be decreasing, that is $dI<0$, then its history is necessarily advancing, which is the main content of the “$I$-Theorem” (Section 1.8.2 of [13]).
@@
The main goal of the present work is to characterize some relevant regimes of the logistic map dynamics from a probabilistic description in terms of the invariant density, from which we calculate FI and associated CR complexity for analyzing dynamical transitions. In addition, we propose a notion of temperature based on the Equipartition Theorem, which allows to provide a macroscopic signature of the map and an analysis of the Frieden’s informational interpretation of the Second Law. The work is organized as follows. In Section II we give the preliminaries. Section III is devoted to the calculation of the Fisher information and the Crámer-Rao complexity, along with our definition of temperature for the logistic map (called map temperature MP). Here we perform numerical simulations of the statistical quantifiers and the MP in function of the parameter of the map and of the time step. Finally, in Section IV, we outline some conclusions and perspectives.
@@ II.1 Probability density functions, discrete maps and invariant density
A probability density function defined over a topological space $\Gamma$ (typically, a subset of $\mathbb{R}^{m}$) is any nonnegative function $p:\Gamma\rightarrow\mathbb{R}_{+}$ such that $\int_{\Gamma}p(x)dx=1$ with $\Gamma$ called the space of events. In an experiment, $\Gamma$ represents the set of all possible outcomes. If these ones are $W$ possible results, then one has a discrete probability distribution expressed by a probability vector $(p_{1},p_{2},\ldots,p_{W})$, being $p_{i}$ the probability that the $i$-th result occurs and the normalization condition reads $\sum_{i=1}^{W}p_{i}=1$.
Given $\Gamma$ and a continuous function $f:\Gamma\rightarrow\Gamma$, we say that the sequence $\{x_{n}\}_{n\in\mathbb{N}_{0}}\subseteq\Gamma$ such that
@@
defines a discrete map. Given $x_{0}\in\Gamma$ as the starting point, the sets $\{x_{n}:n\in\mathbb{N}_{0}\}$ are called orbits, with $x_{\infty}=\lim_{n\rightarrow\infty}x_{n}$ standing for the limit of an orbit. From the physical viewpoint, a discrete map models a system whose dynamics is given by iterating (2), where each iteration corresponds to a time step. If the system is initially in a state $x_{0}$, then $x_{n}$ represents the state after $n$ time steps. When there exists an element $x^{*}$ such that $f(x^{*})=x^{*}$, we say that $x^{*}$ is a fixed point of $f$, which can be physically interpreted as a stationary state of the system. The dynamics of a discrete map can be characterized in terms of probability density functions as follows. The Frobenius–Perron operator $P:\mathbb{L}^{1}(\Gamma)\rightarrow\mathbb{L}^{1}(\Gamma)$ associated with the map (2) is given by [9, 10, 11]
@@
is called the invariant density of the map, constituting a invariant measure [11] that physically represents the state of the system in the asymptotic limit of large iterations [9]. Numerically, the invariant density $\rho(x)$ can be constructed by dividing $\Gamma$ into $W$ intervals and defining
@@
with $1_{[\frac{i-1}{W},\frac{i}{W})}(x)$ standing for the characteristic function of the interval $\frac{i-1}{W},\frac{i}{W})$. It is worth to be noted that in the limit $N\rightarrow\infty$ the invariant density $\rho(x)$ does not depend on the starting point $x_{0}$. Recently, numerical invariant density has been employed to define a statistical distance in discrete maps [18], which resembles the Wooter’s distance [19] for distinguish quantum states.
@@ II.2 Logistic map
where $\mu\in(0,4]$ is the external parameter and $x_{0}\in\Gamma=[0,1]$ the initial condition. The case $\mu=0$ is not of interest since we have $x_{n+1}=0$ for all $n\in\mathbb{N}_{0}$. We review some relevant regimes that we will employ for characterizing them from the formalism presented. When $0<\mu\leq 3$ all the orbits are convergent and we have that $x_{\infty}=0$ for $0<\mu<1$ and $x_{\infty}=\frac{\mu-1}{\mu}$ for $1\leq\mu\leq 3$ with different rates of convergence. For $0<\mu<3$ the orbits rapidly approach to the asymptotic limit with a linear convergence for $2\leq\mu<3$ and less than linear for $\mu=3$. When $3\leq\mu\leq 3.44949$ the orbits oscillate between two values and with $3.44949<\mu\leq 3.56995$ almost all the orbits oscillate between four values. The value $\mu=3.56995$ represents the beginning of the chaotic behavior (onset of chaos) where orbit oscillations of finite period are not observed. The region $3.56995\leq\mu\leq 3.82843$ represents the called Pameau-Maneville scenario [20], with the orbits manifesting a periodic laminar phase interrupted by bursts of aperiodic behavior. For $\mu=4$ the fully chaotic behavior emerges with all the orbits dense in $[0,1]$ provided with a mixing dynamics. Interestingly, when $\mu=2$ and $\mu=4$ there exist exact solutions for the orbits, given by
@@ II.3 Fisher information and Crámer-Rao complexity
The discrete version of the Fisher information (FI) (1) is given in terms of the expression
@@
for a given discrete probability density $\{p_{i}\}$, $p_{i}\geq 0$, normalized as $\sum p_{i}=1$. Associated to (11), we define the Crámer-Rao complexity
@@
with $\sigma^{2}$ standing for the variance of the probability density $\{p_{i}\}$, that is
@@
and $z_{i}\in[\frac{i-1}{W},\frac{i}{W})$ with $p=(p_{1},p_{2},\ldots,p_{W})$ a discrete probability density. As mentioned in the Introduction, the Crámer-Rao complexity is an important statistical quantifier complementary to the Fisher information. The Crámer-Rao bound theorem [13] states that the CR complexity is low bounded by 1, i.e. $C[p]\geq 1$, being maximized for Gaussian probability densities. Here, we only focus on FI and the corresponding CR complexity, being other information measures and their associated complexities [21] out of the scope of the present work.
@@ II.4 Equipartition Theorem
The Equipartition Theorem (ET) establishes a relation between a system’s temperature and its constituent particles total kinetic energy average, expressed in the general form by
@@
where the brackets mean an ensemble average obtained from the Liouville equilibrium density in the asymptotic limit $t\rightarrow\infty$, $H$ is the Hamiltonian of the system, $k_{B}$ the Boltzmann constant, $T$ its temperature and $q_{n}$ denotes the $n$-th coordinate of the system in phase space. For the special case of a system composed by non-interacting particles of mass $m$, the average energy per particle $E_{\textrm{particle}}=\langle H_{\textrm{particle}}\rangle$ is given by the Hamiltonian $H_{particle}=(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})/2m$, so applying (14) for $q_{i}=p_{i}$ ($i=x,y,z$) it follows
@@
from which we can see that the temperature of the system is proportional to the average energy per particle. It is worth noting that some hypotheses are needed for the validity of the ET. Typically, it holds for ergodic systems in thermal equilibrium with a density distribution in phase space equiprobable for all the states. For our proposal, it is convenient to make (15) dimensionless, so considering the one-dimensional case, from (15) we can deduce the dimensionless temperature $T/T_{0}$ as
@@
where $T_{0}$ and $v_{0}$ denote two arbitrary characteristics temperature and velocity for the system.
II.5 Frieden’s informational interpretation of the Second Law of thermodynamics
In order to be compatible the arrow of time with the Fisher information, in Eq. (1.30) of [13] it is has been postulated
$\frac{dI}{dt}\leq 0,$ (17)
which is equivalent to the H-Theorem $\frac{dH}{dt}\geq 0$ for linear Fokker-Planck equations. Physically, equation (17) says that if the Fisher information level of a system is observed to be decreasing, $dI<0$, then its history is necessarily advancing, $dt>0$ (I-Theorem, Section 1.8.2 of [13]). In what follows, equation (17) will be referred as Frieden’s informational interpretation of the Second Law.
@@ III Statistical quantifiers and temperature for logistic map
In this section, we calculate the FI and its corresponding CR complexity for a set of values of the parameter $\mu$ of the logistic map. After that, we analyze the temporal evolution of the FI from the Frieden’s interpretation of Second Law of Thermodynamics for some representative values of $\mu$. Finally, by employing the ET, an associated temperature for the logistic map is obtained along with the FI and the CR complexity as a function of the temperature.
III.1 FI and CR complexity for the logistic map as functions of the map parameter
For the numerical calculations of FI and CR complexity, we employed $N=10^{6}$ steps and $W=10^{4}$ bins, thus guaranteeing $N\gg W$. By increasing $N$, the results remained the same due to the convergence of the invariant density. In Figs. 1 and 2, we illustrate the FI and the CR complexity as functions of the parameter $\mu$.
@@ III.1 FI and CR complexity for the logistic map as functions of the map parameter
From Fig. 2, we see that FI presents an absolute maximum value for $\mu=1$, which corresponds to regular behavior with orbits converging to $x_{\infty}=0$. For $0\leq\mu<1$ and $1\leq\mu<3.4$ the FI behaves approximately constant, from which it is not possible to distinguish the dynamical transitions for these values. From $\mu=3.4$ the FI decreases up to the region $3.8<\mu<3.9$ where a local maximum occurs. FI in the fully chaotic regime $\mu=4$ has the same value as for $3.7\leq\mu\leq 3.8$. By contrast, from Fig. 2 we see that the CR complexity presents a variation in the region $3\leq\mu\leq 4$, which includes the onset of chaos zone from $\mu=3.56995$. A peak is observed between $\mu=3.8$ and $\mu=4$, thus manifesting a detection of the periodic laminar phase with aperiodic behavior characteristic of the Pameau-Maneville scenario, that is a maximal complex behavior. In Fig. 3, we see how the CR complexity takes high values while the FI is vanishingly small and vice-versa.
@@
III.2 Logistic map FI and CR time evolution
With the aim of testing Frieden’s informational interpretation of the Second Law (17), now we analyze the FI and the CR complexity as functions of the time steps. The results are exhibited in Figs. 4 and 5 for regular $\mu=1$, onset of chaos $\mu=3.56995$ and fully chaotic $\mu=4$ regimes. For both FI and CR complexity, we can see that after a short transient, a monotonic behavior is manifested, except for the regular case $\mu=1$, and with appreciable fluctuations for the fully chaotic regime $\mu=4$. When $\mu=1$, we observe that FI is decreasing only by short periods of time, after which it grows abruptly. Hence, the behavior of the FI in the regular regime $\mu=1$ represents a violation of the Frieden’s interpretation of the Second Law $dI/dt\leq 0$ (17). On the other hand, the CR complexity behaves monotonically decreasing for $\mu=1$, as we can see from upper panel of Fig. 5.
Figure 4: Fisher information versus time step of the logistic map for the parameter values $\mu=1$ (upper), $\mu=3.56695$ (center) and $\mu=4$ (bottom) with $N=10^{6}$ number of steps and $W=10^{4}$ bins.
Figure 5: Crámer-Rao complexity versus time step of the logistic map for the parameter values $\mu=1$ (upper), $\mu=3.56695$ (center) and $\mu=4$ (bottom) with $N=10^{6}$ number of steps and $W=10^{4}$ bins.
III.3 A temperature associated to the logistic map
Considering that $v_{i}=x_{i+i}-x_{i}$ represents a dimensionless average velocity for the logistic map between the $i$-th and $(i+1)$-th steps, we apply the relationship (16) with $mv_{0}^{2}/(k_{B}T_{0})=1$ and propose the map temperature (MP) at $N$-th step, denoted by $T(\mu,N,M)$ for the logistic map as follows: